Q5- (5marks) K is normally distributed with a mean of 80 and a 6) The weigh of students in AURA standard deviation of 7. If one students is selcted randomly a- What is the probability of student wight is more than 90 b- What is the probabilit of student wight is between 73 and 87
Answer #1

Consider x = The weight of student in AURAK

x\sim N(\mu =80,\sigma =7)

a) Probability of student weight is more than 90

P[x>90]

=P[z>\frac{90-80}{7}]

=P[Z>1.43]

=1-P[z\leq 1.43]

=1-0.9236...........................by using normal probability table.

=0.0764

b) The probability of student weight is between 73 and 87

=P\left [ \frac{73-80}{7}<z<\frac{87-80}{7} \right ]

=P[-1<z<1]

=P[z\leq 1]-P[z\leq -1]

=0.8413-0.1587.....................by using normal probability table.

=0.6826